OpenStudy (anonymous):

Please Help I will fan and give medals!!!:) Write the quadratic variation equation for the relationship: y varies directly with x2 and y = 36 when x = 3. A.y = 12x B.y = ¼x2 C.y = 4x2 D.y = x2 + 25 @ranga

OpenStudy (ranga):

"y varies directly with x^2" implies y = k * x^2 ---- (1) When x = 3, y = 36 Plug the values into equation (1) and solve for k.

OpenStudy (anonymous):

36=k*9

OpenStudy (anonymous):

then I should divide each side bye 9?

OpenStudy (ranga):

Yes. Divide each side by 9 to solve for k.

OpenStudy (anonymous):

so 4=k?

OpenStudy (ranga):

yes, k = 4. Put k in the equation y = k * x^2 and see which answer choice represents the equation.

OpenStudy (ranga):

Instead of writing x2, I prefer to write it as x^2 (meaning x raised to 2 to indicate that 2 is an exponent).

OpenStudy (anonymous):

36=4*9

OpenStudy (ranga):

You solve for k and got k = 4. Put just k = 4 in the equation y = k * x^2 y = 4x^2 is the equation we are after. Which answer choice has y = 4x^2 (also written as y = 4x2)

OpenStudy (anonymous):

oh duh c.:)

OpenStudy (ranga):

Yep! :)

OpenStudy (anonymous):

Thanks again!

OpenStudy (ranga):

You are welcome.